OpenStudy (rayman5766):

What value of z makes the statement true? –(z + (–17)) = 12 + 17 A. –46 B. –12 C. 12 D. 46

OpenStudy (lieutenantgeneral):

z = -12

OpenStudy (rayman5766):

I knew it

OpenStudy (rayman5766):

Which equation demonstrates the associative property? A. –(33 + 4) = –33 + (–4) B. (–16 + 9) + (–19) = –16 + (9 + (–19)) C. (–5 + 11) + 3 = (11 + (–5)) + 3 D. 2 + (–24) = –24 + 2